if u=q1 q2 and p1 = 2 Naira per kg, p2 = 5 naira per kg and y = p1 q1 + p2 q2.
calculate the quantities of q1 and q2 purchased by the consumer assuming that he was able to maximise his satisfaction if y=200 naira
1
Expert's answer
2021-09-26T20:16:17-0400

given Information

"U = q_1q_2"

P1 = Price of good 1 = 200

P2 = Price of good 2 = 5

 

Budget Constraint is given as 

"p_1 q_1 + p_2 q_2 = 200"

Maximum Utility is achieved where the slope of IC = Sloe if budget constraint.

i.e. "MRS = \frac{P_1}{p_2}"

 

Slope of IC curve

"MRS =\frac{ MU_{q_1}}{MU_{q_2}} \\MU_{q_1} = \frac{∂u}{∂q_1} = q_2\\ MU_{q_2} = \frac{∂u}{∂q_2} = q_1\\MRS = \frac{q_2}{q_1}"

Now Put "MRS = \frac{p_1}{p_2}" , we


"\frac{q_2}{q_1} =\frac{200}{5}\\\frac{ q_2}{q_1} =40\\q_2 = 40q_1\\Now\space Put \space q_2 = 40q_1 \space in\space Budget\space Costraint,\\ we\space get\\p_1 q_1 + p_2 q_2 = 200\\200\times q_1 + 5\times (40q_1) = 200\\200q_1 + 200q_1 = 200\\400q_1 = 200\\q_1 = \frac{1}{2}\\"

Now Put "q_1 = \frac{1}{2}" in budget constraint to calculate the quantity of good 2

"p_1 q_1 + p_2 q_2 = 200\\200\times \frac{1}{2} + 5\times q_2 = 200\\100 + 5q_2 = 200\\5q_2 = 100\\q_2 = 20"


Therefore utility is Maximize when a half unit of good 1 is purchased and 20 units of good 2.